# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.2. Direct limit systems at $\kappa_{i}$
Work in $\mathcal{W}$ and compare $\mathcal{P}^{1}=Lp_{\omega}^{\Sigma_{0},\infty}(\mathcal{W}|\delta_{1})=Lp_{\omega}^{\Sigma,\infty}(\mathcal{W}|\delta_{1})$ via the strategy $\Sigma$ against the construction of $\mathcal{Q}$. Stationarity and universality of fully backgrounded constructions (see [12, Lemmas 2.11 and 2.13]) imply that $\mathcal{P}^{1}$ iterates to some $\mathcal{P}^{*}$ with $\mathcal{P}^{*}\unlhd\mathcal{Q}$. Moreover, by universality, the iteration from $\mathcal{P}^{1}$ to $\mathcal{P}^{*}$ is non-dropping and, by choice of the strategy $\Sigma$, fullness preserving. Write $i\colon\mathcal{P}^{1}\rightarrow\mathcal{P}^{*}$ for the iteration embedding and $\Sigma_{\mathcal{P}^{*}}$ for the corresponding tail strategy of $\mathcal{P}^{*}$. Let $\delta^{*}=i(\delta_{1})$. Then, by fullness preservation, $\mathcal{P}^{*}|((\delta^{*})^{+})^{\mathcal{P}^{*}}=Lp^{\Sigma_{\mathcal{P}^{*}},\infty}(\mathcal{P}^{*}|\delta^{*})$. As $\Sigma$ has branch condensation, it is by [12, Lemma 2.15] absorbed by the fully backgrounded construction resulting in $\mathcal{Q}$. Therefore, in $\mathcal{W}$,
@@
In fact, the argument we just gave shows that there is a club of ordinals $\delta^{*}$ such that $Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*})\vDash\text{``$\delta^{*}$ is Woodin.''}$.” So we can pick an $\eta^{*}$ with $\mathcal{L}^{1}_{\eta^{*}}\cap\operatorname{Ord}=\delta^{*}$ such that $\delta^{*}$ is a strong cutpoint cardinal in $\mathcal{W}$ and $Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*})\vDash\text{``$\delta^{*}$ is Woodin.''}$.”
@@
If (inductively) $\kappa_{0}$ is strong in $\mathcal{M}$, then $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{M}_{\infty}^{0}}$-iterate of $\mathcal{M}_{\infty}^{0}$ in $\mathcal{M}$ and if $\Sigma_{\mathcal{R}_{0}}$ denotes the corresponding tail strategy, $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}\in\mathcal{M}$.
@@
Recall that $\mathcal{R}_{0}$ is constructed in $\mathcal{M}|\delta_{2}$. As we inductively assume that $\kappa_{0}$ is fully strong in $\mathcal{M}$, we have $\Sigma_{\mathcal{M}_{\infty}^{0}}\upharpoonright\mathcal{M}\in\mathcal{M}$ by Lemma 5.11 below (see also the comment after Lemma 5.5). Working in $\mathcal{M}$, compare $\mathcal{M}_{\infty}^{0}$ (via $\Sigma_{\mathcal{M}_{\infty}^{0}}$) against the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ used to define $\mathcal{R}_{0}=\mathcal{L}_{\gamma}$. By stationarity and universality of fully backgrounded constructions (cf. [12, Lemmas 2.11 and 2.13]), $\mathcal{M}_{\infty}^{0}$ iterates to a level of the construction of $\mathcal{R}_{0}$. As $\mathcal{R}_{0}$ is the level where the construction reaches the first full Woodin cardinal, $\mathcal{M}_{\infty}^{0}$ in fact iterates to $\mathcal{R}_{0}$, as desired. This also immediately gives $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}\in\mathcal{M}$. ∎
@@
If (inductively) $\kappa_{0}$ is strong in $\mathcal{M}$, then the function
@@
for $X\in\mathcal{M}|\kappa_{1}$ with $\mathcal{R}_{0}\in X$ is definable over $\mathcal{M}|\delta_{2}$.
@@
The proof is analogous to the proof of Lemma 5.1, replacing fully backgrounded constructions by fully backgrounded constructions relative to $\Sigma_{\mathcal{R}_{0}}$ and using the comparison techniques from [12]. $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}\in\mathcal{M}$ by Lemma 5.7. ∎
For the next level $i=2$, still working in $\mathcal{M}$, let $\delta^{*}\leq\kappa_{2}$ be a $\kappa_{2}$-weak cutpoint and suppose inductively that $\kappa_{0}$ and $\kappa_{1}$ are strong in $\mathcal{M}$. Then we obtain $\mathcal{N}_{\delta^{*}}^{+}$ as follows: Start by performing a fully backgrounded construction above $\delta^{*}$ inside $\mathcal{M}|\delta_{3}$ until the level $\mathcal{R}_{0}$ where it reaches the first full Woodin cardinal. As in Lemma 5.7 we get that $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{M}_{\infty}^{0}}$-iterate of $\mathcal{M}_{\infty}^{0}$ and $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}\in\mathcal{M}$. Therefore, we can add $\Sigma_{\mathcal{R}_{0}}$ to $\mathcal{R}_{0}$ without projecting across $\mathcal{R}_{0}\cap\operatorname{Ord}$. More precisely, if it exists, we let $\mathcal{R}_{1}$ be the result of a fully backgrounded construction relative to $\Sigma_{\mathcal{R}_{0}}$ on top of $\mathcal{R}_{0}$ in $\mathcal{M}|\delta_{3}$ up to a level $\mathcal{R}_{1}$ where the construction reaches the second full Woodin cardinal. That means the construction will not project across $\mathcal{R}_{0}\cap\operatorname{Ord}$ and
@@
All extenders in $\mathcal{M}$ that serve as backgrounds for extenders in $\mathcal{R}_{0}$ or $\mathcal{R}_{1}$ are themselves backgrounded by extenders in $\mathcal{W}$. This follows from the fact that the fully backgrounded constructions giving rise to $\mathcal{R}_{0}$ and $\mathcal{R}_{1}$ are performed in $\mathcal{M}|\delta_{3}$, only use extenders above $\delta^{*}>\delta_{2}$, and stop before $\kappa_{2}$ by Lemma 5.6.
We will show in Lemma 5.9 that $\mathcal{R}_{1}$ is a $\Sigma_{\mathcal{M}_{\infty}^{1}}$-iterate of $\mathcal{M}_{\infty}^{1}$ and, if $\Sigma_{\mathcal{R}_{1}}$ denotes the corresponding tail strategy of $\Sigma_{\mathcal{M}_{\infty}^{1}}$ (and ultimately of $\Sigma_{1}$), $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}\in\mathcal{M}$ (as we are inductively supposing that $\kappa_{0}$ and $\kappa_{1}$ are fully strong in $\mathcal{M}$). Moreover, as strategies with branch condensation are absorbed (see [12, Lemma 2.15]), $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}$ agrees with the strategy $\mathcal{R}_{1}$ inherits from $\mathcal{M}$ (and ultimately from $\mathcal{W}$) as the result of a fully backgrounded construction. So we can let $\mathcal{R}_{2}$ be the result of adding $\Sigma_{\mathcal{R}_{1}}$ to $\mathcal{R}_{1}$. That means, we again construct further relative to $\Sigma_{\mathcal{R}_{1}}$ and, if it exists, obtain $\mathcal{R}_{2}$ as a $\Sigma_{\mathcal{R}_{1}}$-premouse above $\mathcal{R}_{1}$ such that the construction does not project across $\mathcal{R}_{1}\cap\operatorname{Ord}$ and
@@
If (inductively) $\kappa_{0}$ and $\kappa_{1}$ are strong in $\mathcal{M}$, then $\mathcal{R}_{1}$ is a $\Sigma_{\mathcal{M}_{\infty}^{1}}$-iterate of $\mathcal{M}_{\infty}^{1}$ in $\mathcal{M}$ and if $\Sigma_{\mathcal{R}_{1}}$ denotes the corresponding tail strategy, then $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}\in\mathcal{M}$.
@@
Suppose inductively that $\mathcal{M}_{\infty}^{1}\in\mathcal{M}$ and using Lemmas 5.5 and 5.11, $\Sigma_{\mathcal{M}_{\infty}^{1}}\upharpoonright\mathcal{M}\in\mathcal{M}$. Working in $\mathcal{M}$, compare $\mathcal{M}_{\infty}^{1}$ (via $\Sigma_{\mathcal{M}_{\infty}^{1}}$) against the construction of $\mathcal{R}_{1}$ (as the result of a fully backgrounded construction in $\mathcal{M}$) using Sargsyans’s comparison method in [12]. $\mathcal{R}_{1}$ is the result of a fully backgrounded construction using extenders with critical point above $\delta^{*}>\delta_{2}>\mathcal{M}_{\infty}^{1}\cap\operatorname{Ord}$, so (by stationarity and universality of fully backgrounded constructions) $\mathcal{R}_{1}$ does not move in the comparison. Hence, $\mathcal{M}_{\infty}^{1}$ iterates to an initial segment of $\mathcal{R}_{1}$ and their strategies line up. By fullness, $\mathcal{M}_{\infty}^{1}$ in fact iterates to $\mathcal{R}_{1}$. It follows immediately that not only $\Sigma_{\mathcal{M}_{\infty}^{1}}\upharpoonright\mathcal{M}\in\mathcal{M}$ but also its tail strategy $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}$ is in $\mathcal{M}$. ∎
Now returning to the general context for an arbitrary $i>0$, suppose that all $\kappa_{j}$ for $j<i$ are strong in $\mathcal{M}$. We have as in the case $i=0$ that each $\mathcal{N}_{\delta^{*}}^{+}$ for a $\kappa_{i}$-weak cutpoint $\delta^{*}$ is $\kappa_{i}$-suitable and we can prove analogues of Lemmas 5.2 and 5.3. As in the case $i=0$, we let $\mathcal{N}^{*}\in\mathcal{M}$ be the result of making $\mathcal{M}|\kappa_{i}$ generically generic over an iterate of $\mathcal{N}_{\delta_{0}^{*}}^{+}$, where $\delta_{0}^{*}$ is the least $\kappa_{i}$-weak cutpoint in $\mathcal{M}|\kappa_{i}$ above $\delta_{i}$. Recall that as $\mathcal{W}$ is a translatable structure, there is a set of term relations $\vec{\tau}^{(i)}\subseteq\mathcal{N}^{*}$ such that whenever $G\subseteq\operatorname{Col}(\omega,|\mathcal{N}^{*}|)$ is generic over $\mathcal{W}$ and $(\tau_{k}^{(i)}\mid k<\omega)$ is a fixed generic enumeration of $\vec{\tau}^{(i)}$, then $(\Sigma_{\mathcal{N}^{*}})^{G}$ is strongly guided by $(\tau_{k}^{(i)}\mid k<\omega)$. For $l\leq\omega$ let $f_{l}^{(i)}\in F(\kappa_{i},\Sigma)$ be the function given by
@@
$\mathcal{I}^{(i)}=\{(\mathcal{Q},l)\mid\mathcal{Q}\text{ is $\kappa_{i}$-suitable, strongly $f^{(i)}_{l}$-iterable, and countable in $\mathcal{M}^{\operatorname{Col}(\omega,{<}\kappa_{i})}$}\}$
@@
Let $\mathcal{F}^{(i)}$ be the system of models $H^{\mathcal{Q}}_{f_{l}^{(i)}}$ indexed by $(\mathcal{Q},l)\in(\mathcal{I}^{(i)},\leq_{\mathcal{I}^{(i)}})$ together with the associated embeddings. Note that $\mathcal{F}^{(i)}$ is direct and internal to $\mathcal{M}$. Let $(\mathcal{M}_{\infty}^{i})^{*}$ be its direct limit with direct limit embedding $\pi_{(\mathcal{Q},l),\infty}\colon H^{\mathcal{Q}}_{f_{l}^{(i)}}\rightarrow(\mathcal{M}_{\infty}^{i})^{*}$. As $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$ in $\mathcal{M}$, it is not hard to see that there are cofinally many $\kappa_{i}$-weak cutpoints below $\kappa_{i}$ and the $\mathcal{N}_{\delta^{*}}^{+}$’s appear cofinally in $\mathcal{F}^{(i)}$. The following lemma can be shown analogous to Lemma 5.4. The second equality in Lemma 5.10 is a consequence of fullness preservation.
@@
$(\mathcal{M}_{\infty}^{i})^{*}=\mathcal{M}_{\infty}^{i}=(Lp_{\omega}^{\Sigma_{(\mathcal{M}_{\infty}^{i})^{-}},\infty})^{\mathcal{W}}(\mathcal{M}_{\infty}^{i,-})$.
This implies that we can define $\mathcal{M}_{\infty}^{i}$ inside $\mathcal{M}$. As in the case $i=0$, $\mathcal{M}_{\infty}^{i}$ is an iterate of $\mathcal{P}^{i}$ and we write $\pi_{\infty}^{i}\colon\mathcal{P}^{i}\rightarrow\mathcal{M}_{\infty}^{i}$ for the iteration embedding in $\mathcal{W}$. We are left with showing that the corresponding tail iteration strategy $\Sigma_{\mathcal{M}_{\infty}^{i}}$ can be internalized to $\mathcal{M}$ as well. As in the previous section for $i=0$, ideas from the proof of the successor case in the proof of [12, Theorem 6.5] (cf. [13, Lemma 2.15]) can be used to argue that $\Sigma_{\mathcal{M}_{\infty}^{i}}$ is in $\mathcal{M}$. However some new ideas are necessary to deal with the non-fully backgrounded extenders in $\mathcal{M}$. We outline the modifications that yield the internalization of $\Sigma_{\mathcal{M}_{\infty}^{i}}$ to $\mathcal{M}$ also in the case $i>0$.

### v3

@@ 5.2. Direct limit systems at $\kappa_{i}$
Work in $\mathcal{W}$ and compare $\mathcal{P}^{1}=Lp_{\omega}^{\Sigma_{0},\infty}(\mathcal{W}|\delta_{1})=Lp_{\omega}^{\Sigma,\infty}(\mathcal{W}|\delta_{1})$ via the strategy $\Sigma$ against the construction of $\mathcal{Q}$. Stationarity and universality of fully backgrounded constructions (see [13, Lemmas 2.11 and 2.13]) imply that $\mathcal{P}^{1}$ iterates to some $\mathcal{P}^{*}$ with $\mathcal{P}^{*}\unlhd\mathcal{Q}$. Moreover, by universality, the iteration from $\mathcal{P}^{1}$ to $\mathcal{P}^{*}$ is non-dropping and, by choice of the strategy $\Sigma$, fullness preserving. Write $i\colon\mathcal{P}^{1}\rightarrow\mathcal{P}^{*}$ for the iteration embedding and $\Sigma_{\mathcal{P}^{*}}$ for the corresponding tail strategy of $\mathcal{P}^{*}$. Let $\delta^{*}=i(\delta_{1})$. Then, by fullness preservation, $\mathcal{P}^{*}|((\delta^{*})^{+})^{\mathcal{P}^{*}}=Lp^{\Sigma_{\mathcal{P}^{*}},\infty}(\mathcal{P}^{*}|\delta^{*})$. As $\Sigma$ has branch condensation, it is by [13, Lemma 2.15] absorbed by the fully backgrounded construction resulting in $\mathcal{Q}$. Therefore, in $\mathcal{W}$,
@@
In fact, the argument we just gave shows that there is an unbounded set of ordinals $\delta^{*}$ such that for some $\mathcal{Q}$, $Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*})\vDash\text{``$\delta^{*}$ is Woodin.''}$.” Pick an $\eta^{*}$ with $\mathcal{L}^{1}_{\eta^{*}}\cap\operatorname{Ord}=\delta^{*}$ such that $Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*})\vDash\text{``$\delta^{*}$ is Woodin.''}$.”
@@
If $\delta^{*}$ is not a strong cutpoint cardinal in $\mathcal{W}$, we can perform the $\mathcal{P}$-construction inside $\operatorname{Ult}(\mathcal{W},F)$, where $F$ is the least extender overlapping $\delta^{*}$ in $\mathcal{W}$ such that $\delta^{*}$ is a strong cutpoint in $\operatorname{Ult}(\mathcal{W},F)$. As before, this yields that $\delta^{*}$ is not Woodin in $(Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*}))^{\operatorname{Ult}(\mathcal{W},F)}=(Lp_{\omega}^{\Sigma,\infty}(\mathcal{Q}|\delta^{*}))^{\mathcal{W}}$. ∎
@@
The next lemma will be applied in the inductive step from $i=0$ to $i=1$. So we will prove it under the corresponding inductive hypothesis.
@@
Suppose $(\mathsf{IH})_{1}$. Then $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{P}_{\infty}^{0}}$-iterate of $\mathcal{P}_{\infty}^{0}$ in $\mathcal{M}^{(1)}|\delta_{2}=\mathcal{M}^{(0)}|\delta_{2}$ and if $\Sigma_{\mathcal{R}_{0}}$ denotes the corresponding tail strategy, $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}^{(1)}|\delta_{2}$ is amenable to $\mathcal{M}^{(1)}|\delta_{2}$, i.e., for any $X\in\mathcal{M}^{(1)}|\delta_{2}$, $\Sigma_{\mathcal{R}_{0}}\upharpoonright X\in\mathcal{M}^{(1)}|\delta_{2}$.
@@
Recall that $\mathcal{R}_{0}$ is constructed in $\mathcal{M}^{(1)}|\delta_{2}$. As we inductively assume in $(\mathsf{IH})_{1}$ that $\mathcal{M}^{(0)}$ is a well-defined proper class model and $\kappa_{0}$ is fully strong in $\mathcal{M}^{(0)}$, we have that $\Sigma_{\mathcal{P}_{\infty}^{0}}\upharpoonright\mathcal{M}^{(0)}|\delta_{2}$ is amenable to $\mathcal{M}^{(0)}|\delta_{2}$ by Lemma 5.12 below (see also the comment after Lemma 5.5). Moreover, $\mathcal{M}^{(0)}|\delta_{2}=\mathcal{M}^{(1)}|\delta_{2}$ since Clause (2c) in Definition 4.3 only adds extenders with critical point $\kappa_{1}$ above stage $\delta_{2}$ of the construction. Working in $\mathcal{M}^{(1)}|\delta_{2}$, compare $\mathcal{P}_{\infty}^{0}$ (via $\Sigma_{\mathcal{P}_{\infty}^{0}}$) against the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ used to define $\mathcal{R}_{0}=\mathcal{L}_{\gamma}$. By stationarity and universality of fully backgrounded constructions (cf. [13, Lemmas 2.11 and 2.13]), $\mathcal{P}_{\infty}^{0}$ iterates to a level of the construction of $\mathcal{R}_{0}$. As $\mathcal{R}_{0}$ is the level where the construction reaches the first full Woodin cardinal, $\mathcal{P}_{\infty}^{0}$ in fact iterates to $\mathcal{R}_{0}$, as desired. This also immediately gives that $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}^{(1)}|\delta_{2}$ is amenable to $\mathcal{M}^{(1)}|\delta_{2}$. ∎
In fact, the proof straightforwardly generalizes to show the following lemma.
@@
Let $j<\omega$ and suppose $(\mathsf{IH})_{j}$, i.e., $\mathcal{M}^{(j)}|\delta_{j+1}$ is well-defined. Then $\mathcal{R}_{0}$ as constructed in $\mathcal{M}^{(j)}|\delta_{j+1}$ above some $\delta^{*}>\delta_{j}$ is a $\Sigma_{\mathcal{P}_{\infty}^{0}}$-iterate of $\mathcal{P}_{\infty}^{0}$ in $\mathcal{M}^{(j)}|\delta_{j+1}$ and if $\Sigma_{\mathcal{R}_{0}}$ denotes the corresponding tail strategy, $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}^{(j)}|\delta_{j+1}$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$, i.e., for any $X\in\mathcal{M}^{(j)}|\delta_{j+1}$, $\Sigma_{\mathcal{R}_{0}}\upharpoonright X\in\mathcal{M}^{(j)}|\delta_{j+1}$.
Lemma 5.9.
Suppose $(\mathsf{IH})_{1}$. Then the function
@@
for $X\in\mathcal{M}^{(1)}|\kappa_{1}$ with $\mathcal{R}_{0}\in X$ is definable over $\mathcal{M}^{(1)}|\delta_{2}$.
@@
The proof is analogous to the proof of Lemma 5.1, replacing fully backgrounded constructions by fully backgrounded constructions relative to $\Sigma_{\mathcal{R}_{0}}$ and using the comparison techniques from [13]. $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}^{(1)}|\delta_{2}$ is amenable to $\mathcal{M}^{(1)}|\delta_{2}$ by Lemma 5.7. ∎
For the next level $i=2$, suppose $(\mathsf{IH})_{2}$, work in $\mathcal{M}^{(2)}|\delta_{3}$, and let $\delta^{*}\leq\kappa_{2}$ be a $\kappa_{2}$-weak cutpoint. Note that by $(\mathsf{IH})_{2}$, $\mathcal{M}^{(2)}|\delta_{3}$ is well-defined and $\mathcal{M}^{(0)}$ and $\mathcal{M}^{(1)}$ are well-defined proper class models. Moreover, $\mathcal{M}^{(2)}|\delta_{3}=\mathcal{M}^{(1)}|\delta_{3}$. We obtain $\mathcal{N}_{\delta^{*}}^{+}$ as follows: Start by performing a fully backgrounded construction above $\delta^{*}$ inside $\mathcal{M}^{(2)}|\delta_{3}$ until the level $\mathcal{R}_{0}$ where it reaches the first full Woodin cardinal. By Lemma 5.8 we get that $\mathcal{R}_{0}$ is a $\Sigma_{\mathcal{P}_{\infty}^{0}}$-iterate of $\mathcal{P}_{\infty}^{0}$ and $\Sigma_{\mathcal{R}_{0}}\upharpoonright\mathcal{M}^{(2)}|\delta_{3}$ is amenable to $\mathcal{M}^{(2)}|\delta_{3}$. Therefore, we can add $\Sigma_{\mathcal{R}_{0}}$ to $\mathcal{R}_{0}$ without projecting across $\mathcal{R}_{0}\cap\operatorname{Ord}$. More precisely, if it exists, we let $\mathcal{R}_{1}$ be the result of a fully backgrounded construction relative to $\Sigma_{\mathcal{R}_{0}}$ on top of $\mathcal{R}_{0}$ in $\mathcal{M}^{(2)}|\delta_{3}$ up to a level $\mathcal{R}_{1}$ where the construction reaches the second full Woodin cardinal. That means the construction will not project across $\mathcal{R}_{0}\cap\operatorname{Ord}$ and
@@
All extenders in $\mathcal{M}^{(2)}|\delta_{3}$ that serve as backgrounds for extenders in $\mathcal{R}_{0}$ or $\mathcal{R}_{1}$ are themselves backgrounded by extenders in $\mathcal{W}$. This follows from the fact that the fully backgrounded constructions giving rise to $\mathcal{R}_{0}$ and $\mathcal{R}_{1}$ are performed in $\mathcal{M}^{(2)}|\delta_{3}$, only use extenders above $\delta^{*}>\delta_{2}$, and stop before $\kappa_{2}$ by Lemma 5.6.
We will show in Lemma 5.10 that $\mathcal{R}_{1}$ is a $\Sigma_{\mathcal{P}_{\infty}^{1}}$-iterate of $\mathcal{P}_{\infty}^{1}$ and, if $\Sigma_{\mathcal{R}_{1}}$ denotes the corresponding tail strategy of $\Sigma_{\mathcal{P}_{\infty}^{1}}$ (and ultimately of $\Sigma_{1}$), $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}^{(2)}|\delta_{3}$ is amenable to $\mathcal{M}^{(2)}|\delta_{3}$ (as we are inductively supposing $(\mathsf{IH})_{2}$). Moreover, as strategies with branch condensation are absorbed (see [13, Lemma 2.15]), $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}^{(2)}|\delta_{3}$ agrees with the strategy $\mathcal{R}_{1}$ inherits from $\mathcal{M}^{(2)}$ (and ultimately from $\mathcal{W}$) as the result of a fully backgrounded construction. So we can let $\mathcal{R}_{2}$ be the result of adding $\Sigma_{\mathcal{R}_{1}}$ to $\mathcal{R}_{1}$. That means, we again construct further relative to $\Sigma_{\mathcal{R}_{1}}$ and, if it exists, obtain $\mathcal{R}_{2}$ as a $\Sigma_{\mathcal{R}_{1}}$-premouse above $\mathcal{R}_{1}$ such that the construction does not project across $\mathcal{R}_{1}\cap\operatorname{Ord}$ and
@@
Let $j<\omega$ and suppose $(\mathsf{IH})_{j}$, i.e., $\mathcal{M}^{(j)}|\delta_{j+1}$ is well-defined. Then $\mathcal{R}_{1}$ as constructed in $\mathcal{M}^{(j)}|\delta_{j+1}$ above some $\delta^{*}>\delta_{j}$ is a $\Sigma_{\mathcal{P}_{\infty}^{1}}$-iterate of $\mathcal{P}_{\infty}^{1}$ in $\mathcal{M}^{(j)}|\delta_{j+1}$ and if $\Sigma_{\mathcal{R}_{1}}$ denotes the corresponding tail strategy, $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}^{(j)}|\delta_{j+1}$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$, i.e., for any $X\in\mathcal{M}^{(j)}|\delta_{j+1}$, $\Sigma_{\mathcal{R}_{1}}\upharpoonright X\in\mathcal{M}^{(j)}|\delta_{j+1}$.
@@
Suppose inductively that $\mathcal{P}_{\infty}^{1}\in\mathcal{M}^{(j)}|\delta_{j+1}$ and using Lemmas 5.5 and 5.12, $\Sigma_{\mathcal{P}_{\infty}^{1}}\upharpoonright\mathcal{M}^{(j)}|\delta_{j+1}$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$. Working in $\mathcal{M}^{(j)}|\delta_{j+1}$, compare $\mathcal{P}_{\infty}^{1}$ (via $\Sigma_{\mathcal{P}_{\infty}^{1}}$) against the construction of $\mathcal{R}_{1}$ (as the result of a fully backgrounded construction in $\mathcal{M}^{(j)}|\delta_{j+1}$) using Sargsyans’s comparison method in [13]. $\mathcal{R}_{1}$ is the result of a fully backgrounded construction using extenders with critical point above $\delta^{*}>\delta_{j}>\mathcal{P}_{\infty}^{1}\cap\operatorname{Ord}$, so (by stationarity and universality of fully backgrounded constructions) $\mathcal{R}_{1}$ does not move in the comparison. Hence, $\mathcal{P}_{\infty}^{1}$ iterates to an initial segment of $\mathcal{R}_{1}$ and their strategies line up. By fullness, $\mathcal{P}_{\infty}^{1}$ in fact iterates to $\mathcal{R}_{1}$. It follows immediately that not only $\Sigma_{\mathcal{P}_{\infty}^{1}}\upharpoonright\mathcal{M}^{(j)}|\delta_{j+1}$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$ but also its tail strategy $\Sigma_{\mathcal{R}_{1}}\upharpoonright\mathcal{M}^{(j)}|\delta_{j+1}$ is amenable to $\mathcal{M}^{(j)}|\delta_{j+1}$. ∎
Now returning to the general context for an arbitrary $i>0$, suppose $(\mathsf{IH})_{i}$. We have as in the case $i=0$ that each $\mathcal{N}_{\delta^{*}}^{+}$ for a $\kappa_{i}$-weak cutpoint $\delta^{*}$ is $\kappa_{i}$-suitable and we can prove analogues of Lemmas 5.2 and 5.3. As in the case $i=0$, we let $\mathcal{N}^{*}\in\mathcal{M}^{(i)}|\delta_{i+1}$ be the result of making $\mathcal{M}^{(i)}|\kappa_{i}$ generically generic over an iterate of $\mathcal{N}_{\delta_{0}^{*}}^{+}$, where $\delta_{0}^{*}$ is the least $\kappa_{i}$-weak cutpoint in $\mathcal{M}^{(i)}|\kappa_{i}$ above $\delta_{i}$. Recall that as $\mathcal{W}$ is a translatable structure, there is a set of term relations $\vec{\tau}^{(i)}\subseteq\mathcal{N}^{*}$ such that whenever $G\subseteq\operatorname{Col}(\omega,|\mathcal{N}^{*}|)$ is generic over $\mathcal{W}$ and $(\tau_{k}^{(i)}\mid k<\omega)$ is a fixed generic enumeration of $\vec{\tau}^{(i)}$, then $(\Sigma_{\mathcal{N}^{*}})^{G}$ is strongly guided by $(\tau_{k}^{(i)}\mid k<\omega)$. For $l\leq\omega$ let $f_{l}^{(i)}\in F(\kappa_{i},\Sigma)$ be the function given by
@@
$\mathcal{I}^{(i)}=\{(\mathcal{Q},l)\mid\mathcal{Q}\text{ is $\kappa_{i}$-suitable, strongly $f^{(i)}_{l}$-iterable, and countable in $(\mathcal{M}^{(i)})^{\operatorname{Col}(\omega,{<}\kappa_{i})}$}\}$
@@
Let $\mathcal{F}^{(i)}$ be the system of models $H^{\mathcal{Q}}_{f_{l}^{(i)}}$ indexed by $(\mathcal{Q},l)\in(\mathcal{I}^{(i)},\leq_{\mathcal{I}^{(i)}})$ together with the associated embeddings. Note that $\mathcal{F}^{(i)}$ is direct and internal to $\mathcal{M}^{(i)}$. Let $(\mathcal{P}_{\infty}^{i})^{*}$ be its direct limit with direct limit embedding $\pi_{(\mathcal{Q},l),\infty}\colon H^{\mathcal{Q}}_{f_{l}^{(i)}}\rightarrow(\mathcal{P}_{\infty}^{i})^{*}$. As $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$ in $\mathcal{M}^{(i)}$, it is not hard to see that there are cofinally many $\kappa_{i}$-weak cutpoints below $\kappa_{i}$ and the $\mathcal{N}_{\delta^{*}}^{+}$’s appear cofinally in $\mathcal{F}^{(i)}$. The following lemma can be shown analogous to Lemma 5.4. The second equality in Lemma 5.11 is a consequence of fullness preservation.
@@
$(\mathcal{P}_{\infty}^{i})^{*}=\mathcal{P}_{\infty}^{i}=(Lp_{\omega}^{\Sigma_{(\mathcal{P}_{\infty}^{i})^{-}},\infty})^{\mathcal{W}}(\mathcal{P}_{\infty}^{i,-})$.
This implies that we can define $\mathcal{P}_{\infty}^{i}$ inside $\mathcal{M}^{(i)}$. As in the case $i=0$, $\mathcal{P}_{\infty}^{i}$ is an iterate of $\mathcal{P}^{i}$ and we write $\pi_{\infty}^{i}\colon\mathcal{P}^{i}\rightarrow\mathcal{P}_{\infty}^{i}$ for the iteration embedding in $\mathcal{W}$. We are left with showing that the corresponding tail iteration strategy $\Sigma_{\mathcal{P}_{\infty}^{i}}$ can be internalized to $\mathcal{M}^{(i)}$ as well. As in the previous section for $i=0$, ideas from the proof of the successor case in the proof of [13, Theorem 6.5] (cf. [14, Lemma 2.15]) can be used to argue that $\Sigma_{\mathcal{P}_{\infty}^{i}}$ is in $\mathcal{M}^{(i)}$. However some new ideas are necessary to deal with the non-fully backgrounded extenders in $\mathcal{M}^{(i)}$. We outline the modifications that yield the internalization of $\Sigma_{\mathcal{P}_{\infty}^{i}}$ to $\mathcal{M}^{(i)}$ also in the case $i>0$.